# Rethinking the Relationship between the Power Law and Hierarchical Structures

### v2

@@ Abstract
Statistical analysis of corpora provides an approach to quantitatively investigate natural languages. This approach has revealed that several power laws consistently emerge across different corpora and languages, suggesting universal mechanisms underlying languages. Particularly, the power-law decay of correlation has been interpreted as evidence for underlying hierarchical structures in syntax, semantics, and discourse. This perspective has also been extended to child speeches and animal signals. However, the argument supporting this interpretation has not been empirically tested in natural languages. To address this problem, the present study examines the validity of the argument for syntactic structures. Specifically, we test whether the statistical properties of parse trees align with the assumptions in the argument. Using English and Japanese corpora, we analyze the mutual information, deviations from probabilistic context-free grammars (PCFGs), and other properties in natural language parse trees, as well as in the PCFG that approximates these parse trees. Our results indicate that the assumptions do not hold for syntactic structures and that it is difficult to apply the proposed argument to child speeches and animal signals, highlighting the need to reconsider the relationship between the power law and hierarchical structures.
@@ 1 Introduction
Natural languages exhibit characteristic statistical properties. Quantifying these properties through corpus-based analyses is crucial for understanding natural languages, providing insights into the mechanisms that give rise to the properties. Previous research has extensively explored this approach, uncovering various power laws, such as Zipf’s law (Zipf, 1949) and Heaps’ law (Heaps, 1978), that consistently appear across different corpora and languages. These findings suggest the existence of universal mechanisms underlying these power laws.
This study focuses on the power-law decay of correlation. Previous studies have measured correlations between phones, characters, words, and other elements using metrics such as correlation functions and mutual information (MI), across various corpora. These studies have consistently found that the correlation decays according to a power-law function of the distance between elements (Li, 1989a; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Tanaka-Ishii and Bunde, 2016; Lin and Tegmark, 2017; Takahashi and Tanaka-Ishii, 2017; Shen, 2019; Takahashi and Tanaka-Ishii, 2019; Sainburg et al., 2019; Mikhaylovskiy and Churilov, 2023; Nakaishi et al., 2024).
@@
Moreover, the power-law decay of correlation has been interpreted as evidence for the linguistic view that languages are organized hierarchically in syntax, semantics, and discourse (Alvarez-Lacalle et al., 2006; Altmann et al., 2012; Lin and Tegmark, 2017). This interpretation has also been extended beyond speeches and written texts by human adults. Previous research has observed similar power-law behaviors in child speeches and in animal signals, arguing that these may also possess underlying hierarchical structures (Sainburg et al., 2019; Sainburg and Gentner, 2021; Sainburg et al., 2022; Howard-Spink et al., 2024; Youngblood, 2024).
This interpretation is justified by Lin and Tegmark (2017). They showed that the decay of correlation follows a power law in sequences generated by probabilistic context-free grammars (PCFGs; Jelinek et al., 1992), which are a simple model for generating hierarchical tree structures. To be more precise, they proved that in PCFGs, the correlation decays exponentially in structures. In addition, they assumed that the sequential distance grows exponentially with the structural distance. These two give rise to the power-law decay of correlation in sequences.
@@
To address this, we examine whether these assumptions hold for hierarchical structures in natural language syntax, i.e., parse trees. Syntactic structures have been well formalized, resulting in the availability of large-scale treebanks. This enables us to perform statistical analyses on treebanks to test these assumptions.
Specifically, we formalize three key questions essential for examining the validity of the assumptions: (i) Does the correlation decay exponentially in syntactic structures?; (ii) Does the sequential distance grow exponentially with the structural distance?; (iii) To what extent do the statistical properties of syntactic structures deviate from those of PCFGs? To answer these questions, we analyze English and Japanese treebanks.
As a result, our findings are negative for all three questions, indicating that the assumptions in the proposed argument do not hold for syntactic structures in natural languages. Therefore, it is necessary to reconsider the relationship between the two established facts, the power-law decay of correlation and the hierarchical syntactic structures in natural languages.
In addition, we perform the same analyses on trees generated by the PCFG that approximates parse trees, to highlight differences between natural language syntactic structures and those generated by the PCFG. These analyses further show that the MI in structures generated by the PCFG does not exhibit exponential decay when the length of generated sequences is comparable to that of single sentences. Although Lin and Tegmark (2017) proved that the decay converges to an exponential form at sufficiently large distances, which their argument relies on, the decay follows a non-exponential form at distances typically observed in natural language parse trees. The parse trees are too small for this asymptotic exponential behavior to emerge. Consequently, it is difficult to apply their argument not only to single sentences by adults but also child speeches and non-human signals, in which the sequential distance is typically as short as in single sentences.
@@
It would therefore be fruitful to investigate the relationship between the power-law behavior and hierarchical structures at the discourse level. Although previous studies have proposed formal frameworks for discourse annotation (Miltsakaki et al., 2004; Zeldes et al., 2025), the size of existing corpora remain too small to examine the validity of the argument. Developing sufficiently large-scale discourse corpora would be an important next step for future research.
Another direction is to explore alternative mechanisms that could give rise to power-law behavior in natural languages.
@@ 2.2 Power Law and Hierarchical Structures
In studies of human languages and animal signals, this power-law behavior has been interpreted as evidence of underlying hierarchical structures. This interpretation is based on the work of Lin and Tegmark (2017), which shows that the decay of correlation follows a power law in sequences generated by PCFGs.
@@
Lin and Tegmark (2017) have proven that in PCFGs, the correlation decays exponentially with the structural distance, $C\sim\exp(-\lambda r_{\text{str}})$, at sufficiently large distances. Note that the decay may follow non-exponential forms at shorter distances. They also assumed that generated trees are balanced, i.e., the trees have neither left- nor right-branching bias, as in Figure 1 (a). Letting the depth of the two nodes from the common ancestor be $d$, the structural distance is roughly $r_{\text{str}}\sim 2d$, and the sequential distance is asymptotically exponential with $d$, $r_{\text{seq}}\sim q^{d}$, where $q$ is the typical number of children of each node. Therefore, the sequential distance grows exponentially with the structural distance, $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$, where $\mu=(\ln q)/2$. Combining $C\sim\exp(-\lambda r_{\text{str}})$ and $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$ leads to the power-law decay of correlation in sequences, $C\sim r_{\text{seq}}^{-\alpha}$, with $\alpha=\lambda/\mu$.
@@
It is worth noting that, in principle, PCFGs are models not only for syntactic structures but also for more general hierarchical structures. Indeed, PCFGs have been applied to represent hierarchical structures in diverse domains (Ellis et al., 2015; Worth and Stepney, 2005; Gilbert and Conklin, 2007; Knudsen and Hein, 1999; Harlow et al., 2012; Li, 1989b, 1991; Tano et al., 2020; Worth and Stepney, 2021).
From this perspective, the argument by Lin and Tegmark (2017) may also apply to other types of structures, such as semantic and discourse ones. For example, entire texts composed of many sentences intuitively appear to have hierarchical structures defined by paragraphs, sections, chapters, and so on. This hierarchical organization may explain why the decay of correlation follows a power law beyond individual sentences, extending across entire texts.
However, it is difficult to empirically test whether the argument holds for multi-sentence texts. Although previous studies have proposed formal frameworks for annotating discourse structures (Miltsakaki et al., 2004; Zeldes et al., 2025), existing discourse corpora are not yet large enough to reliably estimate the statistical quantities introduced in Section 3, which we use to examine the validity of the assumptions in Lin and Tegmark (2017). Reliable estimation of these quantities requires substantial amounts of data.
@@ 2.3 Rethinking Previous Research
To test whether the argument by Lin and Tegmark (2017) holds for syntactic structures in natural languages, several important questions must be addressed.
@@
Particularly, the second one appears to contradict the well-known fact that syntactic structures typically exhibit a left- or right-branching bias (Hawkins, 1990; Dryer, 1992) (for instance, see Figure 2). In such cases, the growth can be slower than exponential. For example, Figure 1 (b) illustrates a case where the growth is linear, $r_{\text{seq}}\sim r_{\text{str}}$.
In addition, Lin and Tegmark (2017) demonstrate their argument using PCFGs as models of hierarchical structures in natural languages. However, the distribution that PCFGs can represent is highly limited. In this model, the distribution of a subtree depends only on its root, a property we refer to as the context-free independence. Previous studies have suggested that relaxing this property is necessary for modeling syntactic structures (Collins, 2003; Johnson et al., 2006; O’Donnell et al., 2009). Therefore, the last question is:
@@
These three questions remain unresolved. First, although extensive research has measured the correlation in texts and speeches, the correlation in the underlying structures have not been quantified. Therefore, the answer to question (i) remains elusive. Second, syntactic structures exhibit branching bias. Previous studies have not quantitatively examined how this bias affects the relationship between structural and sequential distances, which is crucial for addressing question (ii). Finally, it is well known that PCFGs cannot adequately model natural language syntax without being augmented to relax the context-free independence. However, to answer question (iii) clearly, we need to systematically quantify the differences in the statistical properties between natural language syntax and PCFGs.
In this study, we perform statistical analyses on parse trees to address these three questions, one by one.
@@ 3.1 Datasets and Preprocessing
We primarily use the Brown Laboratory for Linguistic Information Processing 1987-89 Corpus Release 1 (BLLIP; Charniak et al., 2000). Its large scale enables highly precise estimation of the statistical quantities.
@@
The number of child nodes differs across nodes in the dataset, making it difficult to align nodes across different parse trees. To address this issue, we binarize the trees using the Tree.chomsky_normal_form method implemented in the NLTK library (Bird and Loper, 2004). We then apply the Tree.collapse_unary method from the same library to remove the unary branches produced by the former method.
In addition, because a larger number of possible categories requires more data for reliable statistical estimation, we replace part-of-speech (POS) and phrasal tags with a smaller set of categories. Specifically, we refer to the Universal Dependencies (de Marneffe et al., 2021), which contains fewer types of tags than the Penn Treebank-style corpus, and group tags that tend to have similar types of dependents and heads. The eleven categories used in this study are listed in Appendix A. We apply the same categories to both POS and phrasal tags, as the argument by Lin and Tegmark (2017) treats leaves and internal nodes in the same manner. An example of a preprocessed parse tree is shown in Figure 2.
The statistical quantities analyzed in the present study are defined for the distribution of trees and estimated from the empirical dataset. Because parse trees depend on the chosen annotation scheme, these quantities also depend on the scheme. Nevertheless, we expect that the main results, such as whether the decay is exponential or power-law, will remain unchanged as long as the annotation scheme is reasonable.
To verify this, we also perform the same analyses under two alternative settings: trees without binarization and the phrasal-tag-only setting, in which we distinguish phrasal tags from POS tags and focus only on the former. Appendix B shows that the results for both settings are qualitatively similar to those presented in the main text, that is, the answers to questions (i), (ii), and (iii) are negative.
@@
We also analyze the Japanese manually-annotated treebank, NPCMJ (National Institute for Japanese Language and Linguistics, 2016). It contains more than 1.4 million tokens and approximately 90,000 sentences. We first omit all of the phonologically null tokens (null pronouns, traces, etc.) and extract sentences of up to 40 words, resulting in approximately 82,000 sentences. Then we binarize all parse trees and systematically rewrite the POS and phrasal tags using a rule-based mapping scheme listed in Appendix A. Since Japanese is a head-final langauge, all trees are binarized in a left-branching direction, in contrast to the right-branching binarization applied to English.
@@ 3.2 Mutual Information
As a metric for correlations, we use the mutual information (MI) as in Lin and Tegmark (2017). The MI between two tags is defined by
@@
A naive estimate can be obtained by substituting the empirical distributions $\hat{P}(x_{0})$, $\hat{P}(x_{1})$, and $\hat{P}(x_{0},x_{1})$ into $P(x_{0})$, $P(x_{1})$, and $P(x_{0},x_{1})$ in the definitions of entropy. In this case, the estimates of entropy and MI are biased due to finite samples (Arora et al., 2022; Li, 1990).
@@
To further support our claim, we also report the results obtained with six alternative entropy estimators Madow (1948); Miller (1955); Zahl (1977); Horvitz and Thompson (1952); Chao and Shen (2003); Wolpert and Wolf (1995); Nemenman et al. (2001) in Appendix C. The results are consistent with those reported in the main text, except for the estimators that were previously reported to exhibit larger errors (Arora et al., 2022).

### v3

@@ Abstract
Statistical analysis of corpora provides an approach to quantitatively investigate natural languages. This approach has revealed that several power laws consistently emerge across different corpora and languages, suggesting universal mechanisms underlying languages. In particular, the power-law decay of correlations has been interpreted as evidence of underlying hierarchical structures in syntax, semantics, and discourse. This perspective has also been extended beyond corpora produced by human adults, including child speech, birdsong, and chimpanzee action sequences. However, the argument supporting this interpretation has not been empirically tested in natural languages. To address this gap, the present study examines the validity of the argument for syntactic structures. Specifically, we test whether the statistical properties of parse trees align with the assumptions in the argument. Using English and Japanese corpora, we analyze the mutual information, deviations from probabilistic context-free grammars (PCFGs), and other properties in natural language parse trees, as well as in the PCFG that approximates these parse trees. Our results indicate that the assumptions do not hold for syntactic structures and that it is difficult to apply the proposed argument not only to sentences by human adults but also to other domains, highlighting the need to reconsider the relationship between the power law and hierarchical structures.
@@ 1 Introduction
Natural languages exhibit characteristic statistical properties. Quantifying these properties through corpus-based analyses is crucial for understanding natural languages. Previous research has extensively explored this approach, uncovering various power laws, such as Zipf’s law (Zipf, 1949) and Heaps’ law (Heaps, 1978), that consistently appear across different corpora and languages. These findings suggest the existence of universal mechanisms underlying these power laws.
The present study focuses on the power-law decay of correlation. Previous studies have measured correlations between elements such as phones (Sainburg et al., 2019), characters (Li, 1989a; Ebeling and Pöschel, 1994; Ebeling and Neiman, 1995; Lin and Tegmark, 2017; Shen, 2019), and words (Tanaka-Ishii and Bunde, 2016; Takahashi and Tanaka-Ishii, 2017, 2019; Mikhaylovskiy and Churilov, 2023; Nakaishi et al., 2024), using metrics such as correlation functions and mutual information (MI), across various corpora. These studies have consistently found that the correlation decays according to a power-law function of the distance between elements.
@@
Moreover, the power-law decay of correlation has been interpreted as evidence for underlying hierarchical structures. In natural languages, previous research has linked this power-law behavior to linguistic hierarchies in syntax, semantics, and discourse (Alvarez-Lacalle et al., 2006; Altmann et al., 2012; Lin and Tegmark, 2017). This interpretation has also been extended beyond speech and written texts by human adults. Similar power-law behavior has been reported in child speech (Sainburg et al., 2022), birdsong (Sainburg et al., 2019; Sainburg and Gentner, 2021; Youngblood, 2024), and chimpanzee action sequences (Howard-Spink et al., 2024), leading to the claim that these may also possess some form of underlying hierarchical structures, which need not correspond to linguistic hierarchies such as syntax and semantics.
This interpretation is justified by Lin and Tegmark (2017). They showed that the decay of correlation follows a power law in sequences generated by probabilistic context-free grammars (PCFGs; Jelinek et al., 1992), which are a simple model for generating hierarchical tree structures. To be more precise, they proved that in PCFGs, the correlation decays exponentially in sufficiently large structures. In addition, they assumed that the sequential distance grows exponentially with the structural distance. These two give rise to the power-law decay of correlation in sequences.
@@
To address this gap, we test whether these assumptions hold for hierarchical structures in natural language syntax, i.e., parse trees. Syntactic structures have been well formalized, resulting in the availability of large-scale treebanks. This enables statistical analyses to test these assumptions.
Specifically, we formalize three key questions essential for examining the validity of the assumptions: (i) Does the correlation decay exponentially in syntactic structures?; (ii) Does the sequential distance grow exponentially with the structural distance?; (iii) Do the statistical properties of syntactic structures deviate significantly from those of PCFGs? To answer these questions, we analyze English and Japanese treebanks.
Our findings are negative for all three questions, indicating that the assumptions in the proposed argument do not hold for syntactic structures in natural languages. Therefore, it is necessary to reconsider the relationship between the two facts, the power-law decay of correlation and the hierarchical syntactic structures in natural languages.
In addition, we perform the same analyses on trees generated by the PCFG that approximates parse trees, to highlight differences between natural language parse trees and trees generated by the PCFG. These analyses further show that trees generated by the PCFG are too small to exhibit the asymptotic behavior assumed in the argument of Lin and Tegmark (2017). This observation raises doubts about the applicability of their argument not only to single sentences by adults but also to other sequences whose lengths are comparable to those of single sentences, such as child speech, birdsong, and chimpanzee action sequences.
@@
It would therefore be fruitful to investigate the relationship between the power-law behavior and hierarchical structures at the discourse level. Although previous studies have proposed formal frameworks for discourse annotation (Mann and Thompson, 1988; Prasad et al., 2014; Zeldes et al., 2025), the size of existing corpora remain too small to examine the validity of the argument. Developing sufficiently large-scale discourse corpora is an important next step for future research.
Another direction is to explore alternative mechanisms that may give rise to power-law behavior in natural languages. One such mechanism is critical phenomena in the statistical physics literature.
@@ 2.2 Power Law and Hierarchical Structures
In studies of human language, this power-law behavior has been interpreted as evidence of underlying hierarchical structures. This interpretation is based on the work of Lin and Tegmark (2017), which shows that the decay of correlation follows a power law in sequences generated by PCFGs.
@@
Lin and Tegmark (2017) have proven that in PCFGs, the correlation decays exponentially with the structural distance, $C\sim\exp(-\lambda r_{\text{str}})$, at sufficiently large distances. Note that the decay may follow non-exponential forms at shorter distances. They also assumed that generated trees are balanced, i.e., the trees have neither left- nor right-branching bias, as in Figure 1 (a). Letting the depth of the two nodes from the common ancestor be $d$, the structural distance is roughly $r_{\text{str}}\sim 2d$, and the sequential distance is exponential with $d$, $r_{\text{seq}}\sim q^{d}$, where $q$ is the typical number of children of each node. Thus, the sequential distance grows exponentially with the structural distance, $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$, where $\mu=(\ln q)/2$. Combining $C\sim\exp(-\lambda r_{\text{str}})$ and $r_{\text{seq}}\sim\exp(\mu r_{\text{str}})$ leads to the power-law decay of correlation in sequences, $C\sim r_{\text{seq}}^{-\alpha}$, with $\alpha=\lambda/\mu$.
@@
It is worth noting that, in principle, PCFGs can be models not only for syntactic structures but also for more general hierarchical structures. Indeed, PCFGs have been applied to represent hierarchical structures in diverse domains (Ellis et al., 2015; Worth and Stepney, 2005; Gilbert and Conklin, 2007; Knudsen and Hein, 1999; Harlow et al., 2012; Li, 1989b, 1991; Tano et al., 2020; Worth and Stepney, 2021).
From this perspective, the argument by Lin and Tegmark (2017) may also apply to other types of linguistic structures, such as semantic and discourse ones. For example, entire texts composed of many sentences intuitively appear to have hierarchical structures defined by paragraphs, sections, chapters, and so on. This hierarchy may explain why the decay of correlation follows a power law beyond individual sentences, at the level of entire texts.
However, it is difficult to empirically test whether the argument holds for multi-sentence texts. Although previous studies have proposed formal frameworks for annotating discourse structures (Mann and Thompson, 1988; Prasad et al., 2014; Zeldes et al., 2025), existing discourse corpora are not yet large enough to reliably estimate the statistical quantities introduced in Section 3, which we use to examine the validity of the assumptions in Lin and Tegmark (2017).
@@ 2.3 Rethinking Previous Research
To test whether the argument by Lin and Tegmark (2017) holds for syntactic structures in natural languages, several questions must be addressed.
@@
In particular, the second one appears to contradict the well-known fact that syntactic structures typically exhibit a left- or right-branching bias (Hawkins, 1990; Dryer, 1992) (for instance, see Figure 2). In such cases, the growth can be slower than exponential. For example, Figure 1 (b) illustrates a case where the growth is linear, $r_{\text{seq}}\sim r_{\text{str}}$.
In addition, Lin and Tegmark (2017) use PCFGs as models of hierarchical structures in natural languages. However, the expressive power of PCFGs is severely limited. In this model, the distribution of a subtree depends only on its root, a property we refer to as the context-free independence. Previous studies have suggested that relaxing this property is necessary for modeling syntactic structures (Collins, 2003; Johnson et al., 2006; O’Donnell et al., 2009). Therefore, the last question is:
@@
These three questions remain unresolved. First, although extensive research has measured correlations in texts and speech, the correlation in the underlying structures have not been quantified. The answer to question (i) remains elusive. Second, previous studies have not quantitatively examined how the branching bias affects the relationship between structural and sequential distances, which is crucial for addressing question (ii). Finally, it is well known that PCFGs cannot adequately model natural language syntax without being augmented to relax the context-free independence. However, to answer question (iii) clearly, we need to systematically quantify the differences in the statistical properties between natural language syntax and PCFGs.
In this study, we perform statistical analyses on parse trees to address these questions, one by one.
@@ 3.1 Datasets and Preprocessing
We primarily use the Brown Laboratory for Linguistic Information Processing 1987-89 Corpus Release 1 (BLLIP; Charniak et al., 2000). Its large scale enables precise statistical analyses.
@@
The number of child nodes differs across nodes in the dataset, making it difficult to align nodes across different parse trees. To address this issue, we binarize the trees by recursively grouping child nodes from right to left, yielding right-branching structures. As the binarization procedure introduces additional unary branches, we collapse all unary branches. The binarization and unary-collapsing steps were implemented using the Tree.chomsky_normal_form and Tree.collapse_unary methods in the NLTK library (Bird and Loper, 2004). To provide intuition about how much these procedures affect tree shapes, we report depth, branching factor, and the proportion of unary nodes of the original trees in Appendix A.
In addition, we reduce the number of categories, since a larger category set requires more data for reliable statistical estimation. In particular, as categories become more fine-grained, correlations decrease and become difficult to distinguish from noise. Coarse-graining categories is crucial for capturing correlations.
Specifically, we replace part-of-speech (POS) and phrasal tags with a smaller set of categories. We refer to the Universal Dependencies (de Marneffe et al., 2021), which contains fewer types of tags than the Penn Treebank-style corpus, and group tags that tend to have similar types of dependents and heads. The eleven categories used in this study are listed in Appendix B. We apply the same categories to both POS and phrasal tags, as Lin and Tegmark (2017) treat leaves and internal nodes in the same manner. An example of a preprocessed parse tree is shown in Figure 2.
The statistical quantities analyzed here are defined for the distribution of trees and estimated from the empirical dataset. Because parse trees depend on the chosen annotation scheme, these quantities also depend on the scheme. Nevertheless, we expect that the main results, such as whether the decay is exponential or power-law, will remain unchanged unless the annotation scheme is too fine-grained to capture correlations.
To verify this expectation, we also perform the same analyses under two alternative settings: the unbinarized setting and the phrasal-tag-only setting, where we distinguish phrasal tags from POS tags and focus only on the former. Appendix C shows that the results for both settings are qualitatively similar to those in the main text, that is, the answers to questions (i), (ii), and (iii) are negative.
Because the argument of Lin and Tegmark (2017) is based on simple PCFGs, the coarse-grained annotation scheme considered here is sufficient to test the validity of their argument. At the same time, statistical analyses under more refined annotation schemes would be valuable not only for assessing the robustness of our results but also from a linguistic perspective. However, since finer-grained annotation schemes require more data for reliable estimation of statistical quantities, analyses under such schemes would require the construction of annotated corpora much larger than BLLIP. We leave this direction for future work.
@@
We also analyze the Japanese manually-annotated treebank, NPCMJ (National Institute for Japanese Language and Linguistics, 2016). It contains more than 1.4 million tokens and approximately 90,000 sentences. We first remove all of the phonologically null tokens (null pronouns, traces, etc.) and extract sentences of up to 40 words, resulting in approximately 82,000 sentences. We then binarize all parse trees and reduce the number of categories by systematically rewriting the POS and phrasal tags using a rule-based mapping scheme listed in Appendix B. Since Japanese is a head-final langauge, all trees are binarized in a left-branching direction, in contrast to the right-branching binarization applied to English.
As with BLLIP, Appendix A reports depth, branching factor, and the proportion of unary nodes for the original trees in NPCMJ, serving as a reference for assessing both the effect of preprocessing and the differences between English and Japanese. In addition, Appendix C reports results for unbinarized trees. The results are qualitatively similar to those for binarized trees, as in BLLIP, supporting the robustness of our conclusions.
@@ 3.2 Mutual Information
As a metric for correlation, we use the mutual information (MI) as in Lin and Tegmark (2017). The MI between two tags is defined by
@@
A naive estimate can be obtained by substituting the empirical distributions $\hat{P}(x_{0})$, $\hat{P}(x_{1})$, and $\hat{P}(x_{0},x_{1})$ into $P(x_{0})$, $P(x_{1})$, and $P(x_{0},x_{1})$ in the definitions of entropy. In this case, the estimates of entropy and MI are biased due to finite samples (Li, 1990; Arora et al., 2022).
@@
To further support our claim, we also report the results obtained with six alternative entropy estimators Madow (1948); Miller (1955); Zahl (1977); Horvitz and Thompson (1952); Chao and Shen (2003); Wolpert and Wolf (1995); Nemenman et al. (2001) in Appendix D. The results are consistent with those reported in the main text, except for the estimators that were previously reported to exhibit larger errors (Arora et al., 2022).